OpenStudy (aaronandyson):

In an experiment,17 g of ice is used to bring down the temp of 40 g of water at 34 Degree C to its freezing point. sp heat cap of water is 4.2. J G K . cal sp latent heat cap of ice

OpenStudy (aaronandyson):

@imqwerty

imqwerty (imqwerty):

ok the temp changes from 34 to 0 so change in temp =34 \[\Delta Q=ms \DelataT\] \[\Delta Q=40 \times 42 \times 34\] this is amt energy released by water=amt of energy received by ice specific latent heat =energy received/mass =40x42x34/17 =3360J kg^-1

imqwerty (imqwerty):

that is\[\Delta Q=ms \Delta T\]

OpenStudy (aaronandyson):

It should be 336 not 3360

OpenStudy (aaronandyson):

@imqwerty @Michele_Laino

OpenStudy (michele_laino):

I think that the procedure of @imqwerty is right!

OpenStudy (michele_laino):

under the hypothesis that all the ice will be melted

OpenStudy (aaronandyson):

I remember my teacher saying we have equate the equations...

OpenStudy (aaronandyson):

We have 2 equations in the question

OpenStudy (aaronandyson):

that's what she said

OpenStudy (aaronandyson):

heat lost = heat gained...??

OpenStudy (michele_laino):

correct! we have a energy transfer

OpenStudy (aaronandyson):

I can't seem to figure out the equations.. >_>

imqwerty (imqwerty):

do you agree that the amount of heat given by water will all be absorbed by the ice?

OpenStudy (aaronandyson):

yes

imqwerty (imqwerty):

so heat energy given by water=heat energy taken by ice

OpenStudy (aaronandyson):

yes

imqwerty (imqwerty):

can you find how much heat will be given by water to go from 34 degree C to freezing point i.e., 0 degree C?

OpenStudy (aaronandyson):

Q = mL + mc*delta T

imqwerty (imqwerty):

what is L? L=latent heat?

OpenStudy (aaronandyson):

yep

imqwerty (imqwerty):

in this case we are just bringing water to 0 degree C and not converting it into ice so no latent heat will be required so\[Q=ms \Delta T\]

OpenStudy (aaronandyson):

The original has a prob

OpenStudy (aaronandyson):

if the final ans is 336 then it should be 4.2 instead of 42

imqwerty (imqwerty):

i think we have to convert those joules to calories so 4.2J~1cal so this way 42J=4.2cal

OpenStudy (michele_laino):

from your data I read \(4.2 \,J/(grams \cdot K)\)

imqwerty (imqwerty):

OHHH goddd!!!!! I READ THAT 42!!!!!!!!!!!

imqwerty (imqwerty):

>.< D: nvm so we have this- \[Q=ms \Delta T\]\[Q=40 \times 4.2 \times 34\]this is amt of energy received by ice so latent heat=energy received/mass \[LatentHeat=\frac{ 40\times 4.2 \times 32 }{ 17 }=336\] thanks @Michele_Laino !! :)

OpenStudy (michele_laino):

:)

OpenStudy (aaronandyson):

Thanks both of you...!

OpenStudy (michele_laino):

:)